OpenStudy (anonymous):

Simplify the following expression. (2x 2y 4)3 Help Please I have my reasons why I don't understand this.

OpenStudy (whpalmer4):

Is that supposed to be \[(2x*2y*4)^3\]?

OpenStudy (anonymous):

Yes.

OpenStudy (whpalmer4):

Okay. Do you know that \[(ab)^n = a^nb^n\]?

OpenStudy (anonymous):

No I don't know that.

OpenStudy (whpalmer4):

Okay. here's how you can think of that. Let's say you have \(3^4\). That means you are multiplying 3 together 4 times: \(3*3*3*3\) With me so far?

OpenStudy (anonymous):

Yes.

OpenStudy (whpalmer4):

Okay, let's say we have \(6^4\): \(6^4= 6*6*6*6\) right?

OpenStudy (anonymous):

Mhm.

OpenStudy (whpalmer4):

But we can also write 6 as 2*3, so we could write \[6^4 = (2*3)^4 = (2*3)*(2*3)*(2*3)*(2*3)\]and that we can rewrite as \[2*2*2*2*3*3*3*3 = 2^4*3^4\]

OpenStudy (anonymous):

Okay.

OpenStudy (whpalmer4):

So, for your problem, you can just take each thing inside the parentheses, raise it to the 3rd power, and then combine them all by multiplying. Try it and tell me what you get.

OpenStudy (anonymous):

list the options?

OpenStudy (anonymous):

6x6y12 8x5y7 8x6y12 8x8y64

OpenStudy (anonymous):

Uhm Probably won't be close . Lol.

OpenStudy (whpalmer4):

Let's work through it. First thing inside the parentheses is a 2. What is \(2^3\)?

OpenStudy (anonymous):

8x6y12?

OpenStudy (whpalmer4):

okay, if there are exponents, write a ^ in front of them.

OpenStudy (whpalmer4):

for example, \(3^4\) would be written 3^4 (even better would be to use the equation editor, but I'll settle for the ^)

OpenStudy (anonymous):

\[8x ^{2}6y ^{12}\]

OpenStudy (whpalmer4):

Okay, so the original is supposed to be \((2x^2y^4)^3\)? You told me it was something else earlier!

OpenStudy (anonymous):

Ughh I'm confused now.

OpenStudy (whpalmer4):

If the original problem, correctly notated, is to simplify \((2x^2y^4)^3\), then \(8x^6y^{12}\) is the correct answer.

OpenStudy (anonymous):

I meant to say the exponent 6, The two was a typo.

OpenStudy (whpalmer4):

\[2^3 = 8\] \[(x^2)^3 = x^{2*3} = x^6\] \[(y^4)^3 = y^{4*3} = y^{12}\] all of them multiplied together is \(8x^6y^{12}\)

OpenStudy (anonymous):

Okay Thank you.

OpenStudy (whpalmer4):

Sorry about the confusion, but it could have been avoided had the problem been clearly stated up front :-) Got another to do? no extra charge :-)

OpenStudy (anonymous):

Actually I do.

OpenStudy (whpalmer4):

great, let's give it a shot.

OpenStudy (anonymous):

Find the product. \[(m - 5)^{2}\]

OpenStudy (whpalmer4):

Okay, you know how to multiply polynomials, I hope? \[(m-5)^2 = (m-5)(m-5)\]

OpenStudy (anonymous):

You hoped wrong. Lol I'm all new to this.

OpenStudy (whpalmer4):

Okay, well, here's the lightning fast lesson in polynomial multiplication :-)

OpenStudy (anonymous):

Okay :b

OpenStudy (whpalmer4):

Do you know the distributive property of multiplication? \(a(b+c) = ab + ac\)

OpenStudy (whpalmer4):

Right? \(3*15 = 45\) \(3*(10+5) = 3*10 + 3*5 = 30 + 15 = 45\)

OpenStudy (whpalmer4):

Agreed?

OpenStudy (anonymous):

Agree.

OpenStudy (whpalmer4):

Okay. that's all there is to it :-) \[(a+b)(c+d) = a(c+d) + b(c+d) = ac + ad + bc + bd\] In other words, each component of the first polynomial gets multiplied by each component of the second polynomial, and then you add up the pieces. Here's an example with just numbers: \[12*25 = (10+2)(20+5) = 10*20 + 10*5 + 2*20 + 2*5 \]\[= 200 + 50 + 40 + 10 = 300\] So, let's try one with some symbols in there: \[(x+1)(x+2) = \] (you tell me)

OpenStudy (anonymous):

Uhmm

OpenStudy (whpalmer4):

Here, I'll break it down a bit more: \[(x+1)(x+2) = x(x+2) + 1(x+2)\]again by the distributive property

OpenStudy (anonymous):

Can you start me off?

OpenStudy (anonymous):

(x+1)(x+2=(1x+2x)+(1x+2) ?? Hopefully.

OpenStudy (whpalmer4):

so now we distribute again: \[x(x+2) + 1(x+2) = x*x + x*2 + 1(x+2) = x*x + x*2 + 1*x + 1*2\] \[x*x = x^2\]so that becomes \[x^2 + 2x + 1x + 2\]and now we combine the like terms: \[x^2+3x+2\]

OpenStudy (anonymous):

\[4x ^{2}+2\] ??

OpenStudy (whpalmer4):

No, you can't add \(x^2 + 3x\) because they have different exponents. Just like 10 + 1 doesn't equal 20...

OpenStudy (anonymous):

Dang it.

OpenStudy (anonymous):

what to do after coming like terms?

OpenStudy (anonymous):

combing*

OpenStudy (anonymous):

I'm trying to learn I really am, It's so confusing I like just numbers .

OpenStudy (whpalmer4):

after you've combined the like terms, you're done here. I have to go, so why don't we try your actual problem here: \[(m-5)(m-5) = m(m-5) - 5(m-5)\] distributive property, now we do that again with each half of the expression on the right: \[m(m-5) - 5(m-5) = m*m + m(-5) - 5*m -5(-5)\]

OpenStudy (whpalmer4):

that gives us \[m^2 - 5m -5m + 25\] (note that -5*-5 is +25, not -25!) now if we go through and combine like terms: \[m^2-5m-5m + 25 = m^2-10m+25\]so \(m^2-10m+25\) is the answer to the problem.

OpenStudy (anonymous):

Your a freaking Genius. And a life saver Thanks.

OpenStudy (whpalmer4):

You might get some value out of watching the Khan Academy videos on this stuff: https://www.khanacademy.org/math/algebra/polynomials The ones I've looked at have generally been quite well done.